Two boys raising a load pull at an angle to each other. If they exert forces of 30 N and 60 N respectively and their to effective pull is at right angle to the direction of the pull of the first body. What is the angle between their arms ? What is the effective pull?

Grade:11

1 Answers

Alonkrita Bose
5 years ago
Here A= 30N , B=60N , C=90^°
Let D be the angle between 2 forces. Then tan C =BsinD/A+B cosD
or tan 90^°=60sinD/30+60cosD 
or 30+60cosD =0
or cosD=-30/60 =-1/2 =cos120^°
or D=120^°
Effective pulling force is 
R= root A^2+B^2+2ABcosD
=root 900+3600+2×30×60 (-1/2)
=30root3N